MICROWAVE STUDIES OF BENZYL ALCOHOL

\underline{KARISSA G. ATTICKS} AND ROBERT K. BOHN, Department of Chemistry, University of Connecticut, Storrs, CT 06269-3060.

At last year’s Columbus Symposium, Nikolaev and Pratt^{a} reported studies of a benzyl alcohol derivative indicating that the C-O bond does not lie in a plane orthogonal to the benzene ring. That report prompted this study on benzyl alcohol, itself. The microwave rotational spectrum of benzyl alcohol (C_{6}H_{5}CH_{2}OH) has been observed at rotational temperatures near 1 K on a pulsed-jet Fourier transform microwave spectrometer. Measurements have been taken in the range 6 to 15 GHz. A preliminary assignment of the spectrum is consistent with a nonorthogonal orientation of the C-O bond with respect to the benzene ring. Analysis of the spectrum is continuing which will determine the torsional orientation of the C-O bond as well as the conformation of the alcoholic hydrogen.

aA. E. Nikolaev and D. W. Pratt 56th Int. Symp. on Molec. Spectrosc. Paper RF03, June 2001.